for f(x)=-3x^2+12x-12, what is a? and what is the vertex? Is the vertex a minimum or maximum?
Answer by Alan3354(69443)

for f(x)=-3x^2+12x-12, what is a?
a = -3
-------
what is the vertex?
The vertex is at x = -b/2a = -12/-6 = 2
f(2) = -12 + 24 - 12 = 0
Vertex at (2,0)
---------------
Is the vertex a minimum or maximum?
Max